REMPD SPECTROSCOPY OF PROTONATED PAHs FROM THE DISCHARGE SOURCE.

\underline{VADYM A. KAPINUS}, Division Of Chemistry And Chemical Engineering, California Institute Of Technology, Pasadena, CA 91125; and GEOFFREY A. BLAKE, Division Of Geological And Planetary Sciences, California Institute Of Technology, Pasadena, CA 91125.

Protonated polycyclic aromatic hydrocarbons (p-PAH) are predicted to have a few isomers due to multiple protonation sites, but this has yet to be verified experimentally. p-PAHs are viable candidates for being the carriers of the diffused interstellar bands. Resonance Enhanced Multiphoton Dissociation (REMPD) spectra of p-PAHs produced in a longitudinal discharge molecular beam source have been investigated, where the REMPD products were detected with reflectron TOF mass spectrometer.